Value of tan855.how to find this value.this type of question will be asked in neet 2019?

Grade:12th pass

1 Answers

Om
5 years ago
Dear student
tan(855)  =   tan ((2)(360)+135)                                                         [ as tan ((n)(360)+m)  =  tan(m)  ]
= tan (135) =  tan (180-45) =  -tan(45)  =  -1                                      [as tan(45)  =  1     ]
I hope this answer helps you